OpenStudy (anonymous):

What is the greatest common factor (GCF) of the numerator and denominator of the rational expression below? 3x-15/x^2-x-20

OpenStudy (anonymous):

\[x^2-x-20=x^2-5x+4x-20=x \left( ? \right)+4(?)\]